OpenStudy (cakt07):

Solve the problem. The formula d = rt gives the distance traveled in time t at rate r. If a bicyclist rides for 2.3 hours and travels 41 miles, what was her average speed? A. 9.43 miles per hour B. 5.6 miles per hour C. 17.83 miles per hour D. 38.7 miles per hour

OpenStudy (anonymous):

ok... the equation to find the distance is distance = time x rate... which is d = rt if we want to find the rate .. we can write it as rate = distance / time which is r = d/t where d = distance = 41 miles t = time = 2.3 hours so... r= 41/2.3 = 17.83 miles per hour :)

OpenStudy (cakt07):

O-o So... C?

OpenStudy (cakt07):

ohhhhhhhhh! okay i see it now!

OpenStudy (cakt07):

thankyou!